Banyan Co.’s common stock currently sells for $47.75 per share. The growth rate is a constant 12.6%, and the company has an expected dividend yield of 6%. The expected long-run dividend payout ratio is 30%, and the expected return on equity (ROE) is 18%. New stock can be sold to the public at the current price, but a flotation cost of 5% would be incurred. What would be the cost of new equity? Round your answer to two decimal places. Do not round your intermediate calculations.
Answer #1

Stock Price = $47.75

Expected Dividend Yield = D1/Stock Price

So,

D1 = 0.06(47.75)

D1 = $2.865

Flotation Cost = 5%

As per Constant Dividend Growth Model,

Stock Price(1 - f) = D1/(r - g)

r = 2.865/47.75(1 - 0.05) + 0.1260

r = 18.92%

So,

Cost of New Equity = 18.92%